Find two consecutive odd integers such that twice the greater is 17 more than the lesser.
x
x+2

Answer by rapaljer(4671)

You are correct to start with x and x+2.
Let x= smaller odd number
x+2= larger odd number

Now, translate the sentence from English into math symbols.
". . . twice the greater is 17 more than the lesser."
2(x+2) = x+17
2x+4 = x + 17

Subtract x from each side:
2x-x+4= x-x + 17
x+4 = 17

Subtract 4 from each side:
x+4-4=17-4
x=13 Smaller
x+2= 15 Larger

Check:
twice the greater is 17 more than the lesser.
2*15= 13+17
It checks!!

If you need help with WORD PROBLEMS, please see my own website. Do a "Bing" or "Google" search for my last name "Rapalje". Look for "Rapalje Homepage" near the top of the search list. Then, near the top of my Homepage, look for "Basic, Intermediate and College Algebra: One Step at a Time." Click on "Basic Algebra" and look in Chapter 1 for several sections on "Word Problems", especially "Section 1.10 Word Problems, Part I". According to my own students, the explanations are MUCH easier to understand than traditional textbooks, and many of the exercises contained in this section are solved IN COLOR in the corresponding "Math in Living Color" sections.

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus